Peter buys a car worth $21,800 by putting a downpayment of 10% and taking a loan for the balance amount. The loan carries an interest rate of 3.99% over a period of 5 years and needs to be paid on a monthly basis. What is the total interest Peter is expected to pay over the life of the loan?

approximately $2,236

approximately $2,091

approximately $2,067

approximately $2,055

Please show work

Thank you
Answer #1

Loan Amount = 0.90(21,800) = $19,620

Calculating Monthly Payment,

Using TVM Calculation,

PMT = [PV = 19,620, FV = 0, N = 60, I = 0.0399/12]

PMT = $361.24

Total Interest Paid = 60(361.24) - 19,620

Total Interest Paid = $2,055